unseen:

One number is 4 more than another, and their sum is 60. What is the smaller number? 26 28 32

Falconmaster:

26 +30 = 56 so its not a

unseen:

@JustSaiyan

Falconmaster:

32+ 36= 68

Falconmaster:

28+ 32 = 60

Falconmaster:

its b

unseen:

thx

Falconmaster:

welcoem